value: 3.75 points Problem 3-8A Part 7 7. Prepare an adjusted trial balance. RED STORM CLEANERS Adjusted Trial Balance December 31, 2018 ccounts Cash Accounts receivable Supplies Equipment Accumulated depreciation Salaries payable Common stock Retained earnings Dividends Service revenue Salaries expense Repairs and maintenance expense Depreciation expense Supplies expense Debit Credit Totals 0
The general ledger of Red Storm Cleaners at January 1, 2018, includes the following account balances: Debits Credits $12,500 Cash Accounts Recelvable Supplies 6,500 2,500 Accumulated Depreciation Salaries Payable Common Stock Retained Earnings $ 6,000 8500 17500 7000 Totals $39,000 $39,000 The following is ฮ summary of the transactions for the year. a. March 12 Provide services to customers, $45,000, of which $19,500 is on account b. May 2 Collect on accounts recelvable, $16,500 c. June 30 Issue shares of common stock in exchange for $5,000 cash. d. August 1 Pay salaries, $24,500 (of which $8,500 is for saleries peyable in 2017) e. September 25 Pay repairs and maintenance expenses, $1,500 t October 19 Purchase equipment for $6,500 cash g. December 30 Pay $1,000 cash dividends to stockholders h. Accrued salaries at year-end amounted to $1,000. Depreciation for the year on the equipment is $3,500. Office supples remaining on hand at the end of the year equal $1.200
Answer #1
Red Storm Cleaners
Adjusted Trial Balance
31-Dec-18
Particulars Debit Balances Credit Balances
Cash $ 16,000.00
Accounts receivables $ 9,500.00
Supplies $ 1,200.00
Equipment $ 24,000.00
Accumulated depreciation $ 9,500.00
Salaries Payable $ 1,000.00
Common stock $ 22,500.00
Retained Earnings $ 7,000.00
Dividends $ 1,000.00
Service revenue $ 45,000.00
Salaries expense $ 17,000.00
Repairs and maintenance expense $ 11,500.00
Depreciation expense $ 3,500.00
Supplies expense $ 1,300.00
Total $ 85,000.00 $ 85,000.00
Working Note Calculation Amount
Cash = 12500+25500+16500+5000-24500-11500-6500-1000 16000
Accounts receivables = 6500+19500-16500 9500
Equipment = 17500+6500 24000
Accumulated depreciation = 6000+3500 9500
Common stock = 17500+5000 22500
Salaries expense = 16000+1000 17000
Red Storm Cleaners
Adjusted Trial Balance
31-Dec-18
Unadjusted Adjustments Adjusted
Particulars Debit Credit Debit Credit Debit Credit
Cash $ 12,500.00 $ 47,000.00 $ 43,500.00 $ 16,000.00
Accounts receivables $ 6,500.00 $ 19,500.00 $ 16,500.00 $ 9,500.00
Supplies $ 2,500.00 $ 1,300.00 $ 1,200.00
Equipment $ 17,500.00 $ 6,500.00 $ 24,000.00
Accumulated depreciation 6000 3500 $ 9,500.00
Salaries Payable 8500 8500 1000 $ 1,000.00
Common stock 17500 5000 $ 22,500.00
Retained Earnings 7000 $ 7,000.00
Dividends $ 1,000.00 $ 1,000.00
Service revenue $ 45,000.00 $ 45,000.00
Salaries expense $ 17,000.00 $ 17,000.00
Repairs and maintenance expense $ 11,500.00 $ 11,500.00
Depreciation expense $ 3,500.00 $ 3,500.00
Supplies expense $ 1,300.00 $ 1,300.00
Total $ 39,000.00 $ 39,000.00 $ 115,800.00 $ 115,800.00 $ 85,000.00 $ 85,000.00